Jim found out that after working for 9 months he had earned 6 days of vacation time. How many days per year does he earn at this rate?
Please help! Thank you.

Answer by checkley77(12844)

6/9=X/12
9X=6*12
9X=72
X=72/9
X=8 DAYS OF VACATIONS WILL BE EARNED PER YEAR.